OpenStudy (cookiemonster2412):

How far could Jennifer run if she ran 36 km/h for 5 min.

OpenStudy (anonymous):

Convert km/h to m/s and minutes to seconds and plug in formula: \[\frac{ 36 }{ 3.6 } = 10_{m/s}\] \[5 \times 60 = 300_{\sec}\] \[x=v \times t\] \[x = 10 \times 300 = 3000_m\]